A Bibliography of Publications of Herbert H. H. Homeier

Herbert H. H. Homeier Universität Regensburg

Institut fuer Physikalische und Theoretische Chemie Universitätsstr. 31 D-93040 Regensburg

Germany

Tel: +49-941-943 4720 FAX: +49-941-943 4719

E-mail: herbert.homeier@na-net.ornl.gov (Internet)

17 November 2023 Version 1.07

Abstract

This bibliography records publications of Herbert H. H. Homeier.

Title word cross-reference

B [HWS92a, HS93, SH90]. J [Hom95]. J [Hom96e, Hom98c, Hom98d]. Π [Hom97c]. Σ [MHS95].

-separation [MHS95]. -Verfahren [Hom97c].

accelerating [Hom92]. Acceleration [Hom96c, Hom96d, Hom97a, Hom97b, Hom97d, Hom97e, Hom98a, Hom98b, Hom94b]. Accelerators [Hom93]. Addition [HWS92b]. algorithm [Hom92]. am [Hom97c]. Analytical [Hom96e]. anharmonischer [Hom97c]. Anwendung [Hom96b]. Application [Hom97e]. Applications [Hom93]. Asymptotically [Hom98a]. Attraction [HS91, HS93]. Avoiding [Hom97a].

B [Hom90, PMN^{+}92, SHE^{+}98]. B-Funktionen [Hom90]. based [Hom96a, Hom97f]. Basis [HS91, HS92b, HS92a, PMN^{+}92, MHS95]. behavior [Hom96e]. Beispiel [Hom97c]. Berechnung [HD95]. Boundary [HS90].

Calculation [MHS95, PMN^{+}92]. calculations [SHE^{+}98]. Center [HS91, HS92a]. characteristic [Hom97f]. Chemie [Hom96b]. Closures [HRK95]. Coefficients [HS96b]. Combination [PMN^{+}92]. Comment [HS96a]. Complex [Hom97b, Hom97d]. computation [Hom94c, HD95]. consistent [Hom94a, Hom98a]. Contrast [Hom82, HK83, KHdC85]. Convergence [Hom93, Hom96c, Hom96d, Hom97a, Hom97b, Hom97d, Hom97e, Hom98a, Hom98b, Hom99, Hom92, Hom94b, Hom96e]. Convergent [Hom97a]. Correlation [Hom96a, Hom97f]. Coulomb [SHW92]. Coupling [HS96b].

Derivation [HWS92b]. Determinantal [Hom95]. dreidimensionalen [Hom86]. Dyson [Hom96c].

effective [Hom97f]. Effects [Hom82, HK83]. Effektive [Hom86]. einem [Hom86]. Electron [SH90, MHS95]. Energy [Hom96a, Hom97f]. English [HD95, Hom96d]. Equation [HRK95, Hom96c]. Estimates [HW95]. Estimators [Hom96a, Hom97f]. Evaluation [HS92b, HWS92a, HS93, HS96a]. Expansions [Hom97d, Hom98b]. Exponential [HS91, HS92b, HS92a, SHW92]. Exponential-Type [HS91, HS92a, HS92b, SHW92]. Extended [Hom97b]. Extension [Hom97d]. extensivity [Hom97f]. Extrapolation [HRK95, Hom98d]. Extrapolationsverfahren [Hom96b].

Feldtheorie [Hom86]. Field [Hom82, HK83, KHdC85]. Field-Ion [Hom82, HK83]. Fourier [HS92a, Hom92, Hom97b, Hom97e, Hom98a]. Function [PMN^{+}92]. Functions [HS90, HS91, HS92b, HS92a, HWS92a, HS93, SH90, Hom94c, SHE^{+}98]. Funktionen [Hom90].

Gaunt [HS96b]. Gaussian [HS96a].

Harmonics [HS96b]. hierarchically [Hom94a]. Hierarchy [Hom98a]. Hierarchy-consistent [Hom98a]. holes [HD95]. ihre [Hom96b]. Improved [HS91, HS92a]. infinite [HS96a]. Integrals [HS91, HS92b, HWS92a, PMN^{+}92, SH90, SHW92, HS93, HS96a, MHS95]. Integraltransformationsmethoden [Hom90]. Integration [HS90]. Interpolation [Hom99]. interval [HS96a]. Inverse [Hom96c]. involving [HS96a]. Ion [Hom82, HK83, KHdC85]. Iterative [HRK95, Hom96c, Hom98a, Hom94a].

Konvergenzverbesserung [Hom96d].

Landau [Hom86]. Landau-Niveau

[Hom86]. Levin [Hom92, HW95, Hom97e]. Levin-type [Hom92, HW95, Hom97e]. line [HD95]. Linear [Hom99]. Linienform [HD95]. Local [Hom82, HK83]. Löcher [HD95]. Logarithmic [Hom99]. Logarithmically [Hom97a]. Lorentzians [HS96a].

matrices [Hom94c]. Matrizenfolgen [Hom96b]. Mehrreihen [Hom97c]. Mehrreihen- [Hom97c]. Method [Hom97d, Hom98d, MHS95]. Methods [HS91, HS92a, Hom97b, PMN^{+}92]. Microscope [Hom82, HK83, KHdC85]. Möbius [HS90, SH90]. Möbius-Type [SH90]. Molecular [SHE^{+}98]. Molekülintegrale [Hom90]. Møller [Hom96d, Hom96a, Hom96d]. Multicenter [PMN^{+}92, MHS95]. Multipolar [Hom98b].

near [HS90]. Niveau [Hom86]. Nonlinear [Hom93, Hom94b]. Nuclear [HS91, HS93]. Numerical [HS90, Hom96e].

One [HS90, HWS92b]. One-range [HWS92b]. optimal [HS96a]. Orbitals [SHW92]. Ornstein [HRK95]. Orthogonal [Hom97d, Hom98b, Hom94b]. Oszillatoren [Hom97c]. Overlap [HS92b, HWS92a].

Peak [HS90]. Perturbation [Hom96a, Hom96d]. Physikalischen [Hom96b]. Plesset [Hom96d, Hom96a, Hom96d]. polynomials [Hom97f]. Potential [HWS92b]. Product [HS92a]. Programs [HWS92a, HS93]. Progress [SHW92]. Properties [HS96b, Hom97e].

Quadrature [HS91, HS92a, SH90, HS96a]. Quadraturverfahren [Hom90].

range [HWS92b]. Real [HS96b]. Relation [HS96b]. Remainder [HW95]. Representations [SHW92, Hom95]. Repulsion [SH90, MHS95]. Resolution [KHdC85].

semi [HS96a]. semi-infinite [HS96a]. separation [MHS95]. Sequence [HW95, Hom97e, Hom98a, Hom94a]. Series [Hom92, Hom96d, Hom97a, Hom97b, Hom97d, Hom97e, Hom98a, Hom94b]. Sets [PMN^{+}92, MHS95]. shape [HD95]. Sharp [HS90]. Simplified [HWS92b]. size [Hom97f]. size-extensivity [Hom97f]. Slater [MHS95]. Slater-type [MHS95]. Solution [HRK95, Hom96c]. Some [Hom93, HS96b]. spectral [HD95]. spektraler [HD95]. Spherical [HS96b]. Stability [Hom98c]. STO [PMN^{+}92]. Störungsreihe [Hom96d]. studies [Hom96e]. Summation [Hom97a].

Their [HS96b]. Theorem [HWS92b]. Theoretischen [Hom96b]. Theory [Hom96a]. Three [HS91]. Three-Center [HS91]. Transform [HS92a]. Transformation [Hom98a, Hom98c, Hom94a, Hom95, Hom96e]. Transformations [HS90, HW95, Hom97e]. Transforming [Hom99]. Two [HS92a, PMN^{+}92]. Two-Center [HS92a]. Type [HS91, HS92a, SH90, HS92b, Hom92, HW95, Hom97e, MHS95, SHW92].

unterste [Hom86]. Using

[HS90, HRK95, PMN^{+}92, MHS95].

Variations [Hom82, HK83]. Various [HRK95]. Vector [HRK95, Hom98d]. Vektor [Hom96b]. Vektor- [Hom96b]. Verfahren [Hom97c].

Yukawa [HWS92b].

Zahlen  [Hom96b].  Zahlen-

[Hom96b]. Zernike [HRK95]. Zufallspotential [Hom86]. Zur [HD95, Hom96d].

References

Homeier:1995:BLS

[HD95] Herbert H. H. Homeier and B. Dick. Zur Berechnung der Linienform spektraler Löcher (English: On the computation of the line shape of spectral holes). Technical Report TC-PC-95-1, Institut für Physikalische und Theoretische Chemie, Universität Regensburg, D-93040 Regensburg, 1995.

                                                  URL http://www.chemie.uni-
                                                  regensburg.de/preprint.html#

TCPC951. Poster CP 6.15, 59. Physikertagung Berlin 1995, Abstract: Verhandl. DPG (VI) 30, 1815 (1995).

Homeier:1983:ELF

[HK83] Herbert H. H. Homeier and David R. Kingham. Effects of local field variations on the contrast of a field-ion microscope. J. Phys. D, 16:L115–L120, 1983.

Homeier:1982:ELF

[Hom82] Herbert H. H. Homeier. Effects of local field variations on the contrast of a field-ion microscope. Certificate of Postgraduate Study Dissertation, Cambridge University (Great Britain), 1982.

Homeier:1986:EFU

[Hom86] Herbert H. H. Homeier. Effektive Feldtheorie für das unterste Landau-Niveau in einem dreidimensionalen Zufallspotential. Diplomarbeit, Universität Heidelberg, 1986.

Homeier:1990:IQM

[Hom90] Herbert H. H. Homeier. Integraltransformationsmethoden und Quadraturverfahren für Molekülintegrale mit B-Funktionen, volume 121 of Theorie und Forschung. S. Roderer Verlag, Regensburg, Germany, 1990. ISBN 3-89073-482-0. x + 309 pp. LCCN QD462 .H66 1990. URL http:

//www.chemie.uni-regensburg.
de/preprint.html#Homeier90.

Also: Doctoral dissertation, Universität Regensburg.

Homeier:1992:LTA

[Hom92] Herbert H. H. Homeier. A Levintype algorithm for accelerating the convergence of Fourier series. Numer. Algo., 3(1–4):245–254, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Homeier:1993:SAN

[Hom93] Herbert H. H. Homeier. Some applications of nonlinear convergence accelerators. Int. J. Quantum Chem., 45(6):545–562, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Homeier:1994:HCI

[Hom94a] Herbert H. H. Homeier. A hierarchically consistent, iterative sequence transformation. Numer. Algo., 8(1):47–81, 1994. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html{#}TCNA941.

Homeier:1994:NCA

[Hom94b] Herbert H. H. Homeier. Nonlinear convergence acceleration for orthogonal series. In Ralf Gruber and Marco Tomassini, editors, Proceedings of the 6th Joint EPS-APS International Conference on Physics Computing, Physics Computing ’94, pages 47–50. European Physical Society, Boite Postale 69, CH-1213 Petit-Lancy, Geneva, Switzerland, Lugano, 1994. ISBN 2-88270-011-

3. URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCNA942.

Homeier:1994:CFM

[Hom94c] Herbert H. H. Homeier. On the computation of functions of matrices. In S. M. Bachrach, W. Hase, D. B. Boyd, H. S. Rzepa, and S. K. Gray, editors, Proceedings of the First Electronic Computational Chemistry Conference, CD-ROM (ARI-CD 101). ARInternet Corporation, 8201 Corporate Drive Landover, MD 20785, USA, 1994.

URL http://www.chemie.uni-
regensburg.de/ECCC/paper.53/
test.html. Paper 53.

Homeier:1995:DRT

[Hom95] Herbert H. H. Homeier. Determinantal representations for the J transformation. Numer. Math., 71(3):275–288, 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCNA945.

Homeier:1996:CEE

[Hom96a] H. H. H. Homeier. Correlation energy estimators based on Møller–Plesset perturbation theory. J. Mol. Struct. (Theochem), 366:161–171, 1996. URL http:

//www.chemie.uni-regensburg.
de/pub/preprint/preprint.html#
TCQM961.

Homeier:1996:EZV

[Hom96b] H. H. H. Homeier. Extrapolationsverfahren für Zahlen-, Vektor- und Matrizenfolgen und ihre Anwendung in der Theoretischen und Physikalischen Chemie. Habilitation thesis, Universität Regensburg, Germany, 1996.

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#Homeier96Hab.

Homeier:1996:CAI

[Hom96c] H. H. H. Homeier. On convergence acceleration for the iterative solution of the inverse Dyson equation. J. Mol. Struct. (Theochem), 368:81–91, 1996.

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCQM954. Pro-

ceedings of the 2nd Electronic Computational Chemistry Conference.

Homeier:1996:KMP

[Hom96d] H. H. H. Homeier. Zur Konvergenzverbesserung der Møller– Plesset Störungsreihe (English: On convergence acceleration of the Møller–Plesset perturbation series). Technical Report Homeier:1996:KMP, Institut für Physikalische und Theoretische Chemie, Universität Regensburg, D-93040 Regensburg, 1996.

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCQM962. Poster

CP 14.77, Frühjahrstagung des Arbeitskreises Festkörperphysik bei der DPG, Regensburg 1996. Abstract: Verhandl. DPG (VI) 31, 2165-2166 (1996).

Homeier:1996:ANS

[Hom96e] Herbert H. H. Homeier. Analytical and numerical studies of the convergence behavior of the J transformation. J. Comput. Appl. Math., 69:81–112, 1996.

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCNA944.

Homeier:1997:CAL

[Hom97a] H. H. H. Homeier. Convergence acceleration of logarithmically convergent series avoiding summation. Technical Report Homeier:1997:CAL, Institut für Physikalische und Theoretische Chemie, Universität Regensburg, D-93040 Regensburg, 1997.

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCNA978. Appl.

Math. Lett., in press.

Homeier:1997:ECSa

[Hom97b] H. H. H. Homeier. Extended complex series methods for the convergence acceleration of Fourier series. Technical Report Homeier:1997:ECSa, Institut für Physikalische und Theoretische Chemie, Universität Regensburg, D-93040 Regensburg, 1997.

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCNA973.  J.

Comput. Phys., submitted.

Homeier:1997:MVB

[Hom97c] H. H. H. Homeier. Mehrreihen-Π-Verfahren am Beispiel anharmonischer Oszillatoren. Technical Report Homeier:1997:MVB, Institut für Physikalische und Theoretische Chemie, Universität Regensburg, D-93040 Regensburg, 1997. URL http:

//www.chemie.uni-regensburg.
de/pub/preprint/preprint.html#

TCQM973. Poster, 33. Symposium für Theoretische Chemie, Köln-Walberberg, Germany.

Homeier:1997:ECSb

[Hom97d] H. H. H. Homeier. On an extension of the complex series method for the convergence acceleration of orthogonal expansions. Technical Report Homeier:1997:ECSb, Institut für Physikalische und Theoretische Chemie, Universität Regensburg, D-93040 Regensburg, 1997.

                                        URL http://www.chemie.uni-
                                        regensburg.de/pub/preprint/
                                        preprint.html#TCNA974. Numer.

Math., submitted.

Homeier:1997:PAL

[Hom97e] H. H. H. Homeier. On properties and the application of Levintype sequence transformations for the convergence acceleration of Fourier series. Technical Report Homeier:1997:PAL, Institut für Physikalische und Theoretische Chemie, Universität Regensburg, D-93040 Regensburg, 1997.

                                        URL http://www.chemie.uni-
                                        regensburg.de/pub/preprint/
                                        preprint.html#TCNA971. Math.

Comp., submitted.

Homeier:1997:SEC

[Hom97f] H. H. H. Homeier. The sizeextensivity of correlation energy estimators based on effective characteristic polynomials. J. Mol. Struct. (Theochem), 419:29–31, 1997. URL http:

                                        //www.chemie.uni-regensburg.
                                        de/pub/preprint/preprint.html#

TCQM963. Proceedings of the 3^{rd} Electronic Computational Chemistry Conference.

Homeier:1998:AHC

[Hom98a] H. H. H. Homeier. An asymptotically hierarchy-consistent iterative sequence transformation for convergence acceleration of Fourier series. Numer. Algo., 18(1):1–30, 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCNA972.

Homeier:1998:CAM

[Hom98b] H. H. H. Homeier. On convergence acceleration of multipolar and orthogonal expansions. Internet J. Chem., 1(Article 28):????,

1998. URL http://www.ijc.com/
articles/1998v1/28/. Proceed-

ings of the 4^{th} Electronic Computational Chemistry Conference.

Homeier:1998:ST

[Hom98c] H. H. H. Homeier. On the stability of the J transformation. Numer. Algo., 17(3–4):223–239, 1998. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCNA976.

Homeier:1998:VEM

[Hom98d] H. H. H. Homeier. The vector J extrapolation method. In Junping Wang, Myron B. Allen, III., Benito M. Chen, and Tarek Mathew, editors, Iterative Methods in Scientific Computation, volume 4 of IMACS Series in Computational and Applied Mathematics, pages 375–380. IMACS, Dept. of Computer Science, Rutgers University, New Brunswick, NJ 08903, USA, 1998. ISSN 1098-870X.

                                        URL http://www.chemie.uni-
                                        regensburg.de/pub/preprint/
                                        preprint.html#TCNA975. Pro-

ceedings of the Third IMACS International Symposium on Iterative Methods in Scientific Computation Jackson Hole, Wyoming, USA - July 9-12, 1997.

Homeier:1999:TLL

[Hom99] H. H. H. Homeier. Transforming logarithmic to linear convergence by interpolation. Appl. Math. Lett., 12:13–17, 1999.

                                        URL http://www.chemie.uni-
                                        regensburg.de/pub/preprint/
                                        preprint.html#TCNA976.

Homeier:1995:ISO

[HRK95] H. H. H. Homeier, S. Rast, and H. Krienke. Iterative solution of the Ornstein–Zernike equation with various closures using vector extrapolation. Comput. Phys. Commun., 92:188–202, 1995.

                                        URL http://www.chemie.uni-
                                        regensburg.de/pub/preprint/
                                        preprint.html{#}TCPC952.

Homeier:1990:NIF

[HS90] Herbert H. H. Homeier and E. Otto Steinborn. Numerical integration of functions with a sharp peak at or near one boundary using Möbius transformations. J. Comput. Phys., 87:61–72, 1990.

Homeier:1991:IQM

[HS91] Herbert H. H. Homeier and E. Otto Steinborn. Improved quadrature methods for threecenter nuclear attraction integrals with exponential-type basis functions. Int. J. Quantum Chem., 39 (4):625–645, April 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Homeier:1992:IQM

[HS92a] Herbert H. H. Homeier and E. Otto Steinborn. Improved quadrature methods for the Fourier transform of a two-center product of exponential-type basis functions. Int. J. Quantum Chem., 41(3):399–411, February 5, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Homeier:1992:EOI

[HS92b] Herbert H. H. Homeier and E. Otto Steinborn. On the evaluation of overlap integrals with exponential-type basis functions. Int. J. Quantum Chem., 42(4): 761–778, May 20, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Homeier:1993:PEN

[HS93] Herbert H. H. Homeier and E. Otto Steinborn. Programs for the evaluation of nuclear attraction integrals with B functions. Comput. Phys. Commun., 77:135– 151, 1993.

Homeier:1996:CGQ

[HS96a] H. H. H. Homeier and E. O. Steinborn. Comment on: A Gaussian quadrature for the optimal evaluation of integrals involving Lorentzians over a semi-infinite interval. Comput. Phys. Commun., 99(1):77–80, 1996.

Homeier:1996:SPC

[HS96b] H. H. H. Homeier and E. O. Steinborn. Some properties of the coupling coefficients of real spherical harmonics and their relation to Gaunt coefficients. J. Mol. Struct. (Theochem), 368:31–37, 1996.

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCQM953. Pro-

ceedings of the 2nd Electronic Computational Chemistry Conference.

Homeier:1995:REL

[HW95] H. H. H. Homeier and E. J. Weniger. On remainder estimates for Levin-type sequence transformations. Comput. Phys. Commun., 92:1–10, 1995. URL http:

//www.chemie.uni-regensburg.
de/pub/preprint/preprint.html#
TCNA951.

Homeier:1992:PEO

[HWS92a] Herbert H. H. Homeier, E. Joachim Weniger, and E. Otto Steinborn. Programs for the evaluation of overlap integrals with B functions. Comput. Phys. Commun., 72:269–287, 1992.

Homeier:1992:SDO

[HWS92b] Herbert H. H. Homeier, E. Joachim Weniger, and E. Otto Steinborn. Simplified derivation of a one-range addition theorem of the Yukawa potential. Int. J. Quantum Chem., 44(3):405–411, September 30, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kingham:1985:RCF

[KHdC85] David R. Kingham, Herbert H. H. Homeier, and Caio M. C. de Castilho. Resolution and contrast of a field ion microscope. Surf. Sci., 152/153:55–61, 1985.

Maslov:1995:CME

[MHS95] I. V. Maslov, Herbert H. H. Homeier, and E. Otto Steinborn. Calculation of multicenter electron repulsion integrals in Slater-type basis sets using the Σ-separation method. Int. J. Quantum Chem., 55(1):9–22, July 5, 1995. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

URL http://www.chemie.uni-
regensburg.de/pub/preprint/
preprint.html#TCMI941.

Perevozchikov:1992:CTM

[PMN^{+}92] V. I. Perevozchikov, I. V. Maslov, A. W. Niukkanen, Herbert H. H. Homeier, and E. Otto Steinborn. On the combination of two methods for the calculation of multicenter integrals using STO and B function basis sets. Int. J. Quantum Chem., 44(1):45–57, August 20, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Steinborn:1990:MTQ

[SH90] E. Otto Steinborn and Herbert H. H. Homeier. Möbius-type quadrature of electron repulsion integrals with B functions. Int. J. Quantum Chem. Symp., 24:349– 363, 1990.

Steinborn:1998:MCB

[SHE^{+}98] E. O. Steinborn, H. H. H. Homeier, I. Ema, R. López, and G. Ramírez. Molecular calculations with B functions. Technical Report Steinborn:1998:MCB, Institut für Physikalische und Theoretische Chemie, Universität Regensburg, D-93040 Regensburg, 1998.

                                        URL http://www.chemie.uni-
                                        regensburg.de/pub/preprint/
                                        preprint.html#TCQM981. Int. J.

Quantum Chem., in press.

Steinborn:1992:RPR

[SHW92] E. Otto Steinborn, Herbert H. H. Homeier, and E. Joachim Weniger. Recent progress on representations for Coulomb integrals of exponential-type orbitals. J. Mol. Struct. (Theochem), 260:207–221, 1992.